# Principled Federated Random Forests for Heterogeneous Data

### v1

@@ 1 Introduction
Random Forests (RFs) (Breiman, 2001) are ensembles of CART decision trees (Breiman et al., 1984) that recursively partition the feature space into regions, referred to as nodes, by selecting splits that maximize an impurity reduction criterion (e.g., variance or Gini). Predictions are then produced by aggregating the outcomes associated with the terminal nodes, commonly called leaves. Despite the deep learning revolution, RFs remain among the state-of-the-art and most widely used predictors for tabular data (Fernández-Delgado et al., 2014; Grinsztajn et al., 2022; Kaggle, 2022; Uddin and Lu, 2024), while also offering practical advantages: they naturally handle both continuous and categorical features, require minimal hyperparameter tuning, and are substantially more computationally efficient than deep learning models (Grinsztajn et al., 2022). The strong empirical performance of RFs has been established in centralized settings, where data are assumed to be pooled together. However, this assumption is often unrealistic in many real-world applications, particularly in healthcare, where data are distributed across multiple institutions and cannot be centralized due to privacy regulations, governance constraints, or competitive considerations (Rieke et al., 2020; Antunes et al., 2022; Nguyen et al., 2022; Xu et al., 2021). Federated learning (FL) addresses this challenge by enabling collaborative model training on decentralized data (Kairouz et al., 2021). In this work, we focus on the standard server–client architecture and on horizontal FL, where clients hold disjoint sets of records described by a common feature space.
@@
A recent survey by Wang and Gai (2024) identifies three key challenges in decentralized training of tree-based models: computing split statistics without leaking sensitive information, limiting communication costs, and handling client heterogeneity.
@@
Although RFs enjoy a substantial theoretical foundation in the centralized setting (Scornet et al., 2015), analogous guarantees are currently lacking for their federated counterparts. To pave the way toward provably reliable federated forests, we study the fundamental question of whether a federated split-selection mechanism can recover, or closely approximate, the greedy split decisions made by centralized CART.
@@ 2.1 Centralized CART/RF Split Selection
The mixture-risk objective (1) is an ideal target. CART and Random Forests approximate it through greedy, node-wise optimization of an impurity criterion (variance for regression; Gini or entropy for classification) on a finite data sample. In the centralized setting, all data are pooled as $\mathcal{D}=\bigcup_{k=1}^{K}\mathcal{D}_{k}$. At a node $\nu\subseteq\mathcal{X}$ of a tree, let $\mathcal{S}_{\nu}\subseteq\mathcal{D}$ be the samples reaching $\nu$. CART selects a split threshold (j,t)(j,t)–that is, a split variable $j\in[d]$ and a cutpoint $t\in\mathbb{R}$–in two steps: (i) candidate generation, where for each feature $j$ the values $\{x_{i,j}\}_{i\in\mathcal{S}_{\nu}}$ are sorted, and candidates $t$ are defined as the midpoints between consecutive distinct values; and (ii) split evaluation, where each candidate $(j,t)$ is scored by its empirical impurity reduction
@@
with $\mathcal{S}_{L}=\{i\in\mathcal{S}_{\nu}:x_{i,j}\leq t\}$, $\mathcal{S}_{R}=\{i\in\mathcal{S}_{\nu}:x_{i,j}>t\}$, $n_{L}=|\mathcal{S}_{L}|$, $n_{R}=|\mathcal{S}_{R}|$, and $n_{\nu}=n_{L}+n_{R}$. CART then selects the threshold $(j,t)$ that maximizes $\Delta I(j,t;\mathcal{S}_{\nu})$.
@@ 2.2 Heterogeneity Scenarios
Figure 1: Graphical models of the heterogeneity regimes. Dashed arrows indicate possible dependencies between the client variable $H$ and the covariates and/or outcome mechanism, depending on the regime.
General model.
In full generality, the combined data admit the decomposition
@@
Assumption 2.2 (Covariate shift setting).
@@
Assumption 2.3 (Outcome shift (client effect) setting).
$P(X\mid H)=P(X)$ while $P(Y|X)$ may depend on $H$.
@@ 3.1 Candidate Split Generation
Theorems 3.1 and 3.2 (proved in Appendix A.3) show that $B$ directly controls the rank resolution of the candidate grid: increasing $B$ makes the federated quantile thresholds closer to the centralized midpoint thresholds in terms of induced child assignments. For example, setting $B=100$ percentiles implies at most $1.5\%$ disagreement in child assignments while requiring only B1B-1 candidates per feature, independent of $n_{\nu}$. These results allow precise control of the trade-off between communication cost and candidate-threshold fidelity to the centralized approach.
Our quantile-based candidates are superior to the fixed-width histograms proposed by Kalloori and Klingler (2022) for three reasons: (i) they adapt to local data density, ensuring critical split points are preserved as well as possible for a given budget $B$; (ii) they are more robust against outliers; and (iii) they mitigate privacy leakage by enforcing a minimum sample fraction of $1/B$ within each bin.
Top-$L$ feature shortlisting for homogeneous data. Under 2.1, local impurity gains concentrate around the pooled gain, so local rankings tend to agree with the global ranking. This enables a communication-efficient preselection step before candidate generation: each client $k$ computes, for each feature $j$, its best local gain
$M_{\nu,k}(j)=\max_{t}\ \Delta I\big(j,t;\mathcal{S}_{\nu,k}\big),$
and sends only its top-$L$ features (with associated gains). The server merges these reports (e.g., taking the union of the local top-$L$ sets) to form a reduced feature set $\mathcal{L}_{\nu}$ on which it performs quantile sketching and split evaluation. A supporting guarantee for retaining the pooled-best feature under 2.1 is given in Appendix A.4.1.
@@ 3.2 Federated Impurity Reduction
Given the candidate set from Section 3.1, the remaining step is to evaluate the impurity reduction (2) for each candidate split $(j,t)$ without exchanging raw data subsets. Our main evaluation rule is based on a novel observation: standard CART impurities admit additive sufficient statistics that can be aggregated across clients to recover the centralized impurity exactly.
Exact centralized evaluation via additive summaries. For standard CART criteria, there exists a deterministic map $\Psi$ and per-client summaries $\mathbf{s}_{\nu,k}$ such that the pooled impurity reduction can be written as
@@
with aggregated statistics $\mathbf{s}_{m}=\sum_{k=1}^{K}\mathbf{s}_{m,k}$ for $m\in\{\nu,L,R\}$. In practice, for a candidate $(j,t)$, each client sends only its left-child summary $\mathbf{s}_{L,k}$; the server aggregates $\mathbf{s}_{L}$, recovers $\mathbf{s}_{R}=\mathbf{s}_{\nu}-\mathbf{s}_{L}$, and evaluates (5). This matches centralized split evaluation exactly, regardless of covariate shift or outcome heterogeneity.
Examples. For $C$-class classification, $\mathbf{s}_{\nu}=(n_{\nu,1},\dots,$ $n_{\nu,C})$, $\Psi_{\mathrm{Gini}}(\mathbf{s}_{\nu})=1-\sum_{c=1}^{C}(n_{\nu,c}/n_{\nu})^{2}$, so each client sends $C$ scalars per candidate. For regression, $\mathbf{s}_{\nu}=(n_{\nu},S_{\nu},Q_{\nu})$ with $S_{\nu}=\sum y_{i}$, $Q_{\nu}=\sum y_{i}^{2},\Psi_{\mathrm{Var}}(\mathbf{s}_{\nu})=Q_{\nu}/n_{\nu}-(S_{\nu}/n_{\nu})^{2}$, so each client sends $3$ scalars.
The following theorem justifies the above strategy by showing that aggregating local impurities generally fails to recover the global impurity, introducing an additional error term.
Theorem 3.3 (General impurity decomposition).
Let $I(\mathcal{S}_{\nu})$ denote the empirical impurity at node $\nu$. Then
$I(\mathcal{S}_{\nu})=\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}I(\mathcal{S}_{\nu,k})+\mathcal{E}(\nu),$ (6)
where $\mathcal{E}(\nu)$ is a heterogeneity (Jensen-gap) term induced by the within-node mixture of client distributions. For standard strictly concave impurity functions (e.g., Gini index, entropy, variance), we have $\mathcal{E}(\nu)\geq 0$, with $\mathcal{E}(\nu)=0$ if and only if
i.e., the within-node empirical outcome means (or class probabilities) coincide across all clients.
The term $\mathcal{E}(\nu)$ explains why split decisions based solely on local objectives can deviate from the centralized CART choice when clients are heterogeneous.
Cheaper approximation based on averaging local gains (AvgImp). While the decomposition above highlights risks in heterogeneous settings, it also suggests a faster alternative for homogeneous regimes (2.1). When $\mathcal{E}(\nu)\approx 0$, the global gain can be approximated by the weighted average of local gains:
This proxy motivates our AvgImp strategy: a conjunction of (i) local Top-$L$ screening—where each client proposes only its best candidates based on local gains—and (ii) an evaluation based on the averaged metric above. This approach drastically reduces communication by avoiding the evaluation of all pooled candidates.
Theorem 3.4 (AvgImp approximation error on homogeneous data).
Under 2.1 and $\mathbb{E}[Y^{2}]<\infty$,
In summary, (5) is our primary evaluation rule (exact centralized gain), while Theorem 3.3Theorem 3.4 motivate AvgImp as a communication-lean approximation that is accurate when data is (close to) homogeneous.
Outcome shift: splitting on $H$ with no extra communication. Under 2.3, the target rule is $f(x,h)=\mathbb{E}[Y\mid X=x,H=h]$, and optimal trees may require splits on $H$. Let $\mathcal{K}_{\nu}=\{k:\ n_{\nu,k}>0\}$ be the set of active clients at node $\nu$. Once the server has collected the per-client node summaries $\{\mathbf{s}_{\nu,k}\}_{k\in\mathcal{K}_{\nu}}$—already needed to evaluate feature-based splits—it can also evaluate splits on $H$ at no additional communication cost, i.e., without requesting any extra messages from clients. Indeed, any partition $\mathcal{K}_{\nu}=\mathcal{K}_{L}\cup\mathcal{K}_{R}$ induces
$\mathbf{s}_{\nu,L}=\sum_{k\in\mathcal{K}_{L}}\mathbf{s}_{\nu,k}\qquad\text{and}\qquad\mathbf{s}_{\nu,R}=\sum_{k\in\mathcal{K}_{R}}\mathbf{s}_{\nu,k},$
and the corresponding gain follows directly from (5) using the same map $\Psi$.
@@ 4 Federated Forests
We now describe how to transform the federated split primitives of Section 3—candidate generation and split evaluation from summaries—into a scalable Random Forest training protocol, which we call FedForest. The key ingredients are: (i) local bootstrap resampling to introduce tree-level randomness while preserving the pooled mixture, (ii) per-node feature subsampling and (optional) Top-$L$ shortlisting to limit the number of feature–threshold pairs evaluated, and (iii) level-wise batching to train many trees in a small number of synchronized communication rounds.
4.1 Bootstrap and Client Subsampling
Client-stratified bootstrap. For each tree, client $k$ draws (with replacement) a bootstrap sample of size $n_{k}$ from $\mathcal{D}_{k}$ and the tree is trained on the virtual union of these resamples across clients. This preserves the pooled mixture weights $\rho_{k}=n_{k}/n$ in every tree. In contrast, a pooled bootstrap can under-represent—or entirely omit—small clients.
Client subsampling. For a given tree (or node), the server can a client subset $\mathcal{K}^{\prime}\subseteq[K]$ to reduce communication. We recommend it mainly as an acceleration knob under approximately homogeneous clients: under covariate shift, it changes the effective mixture distribution and may alter split decisions, and under outcome shift it can also omit informative values of $H$.
4.2 Parallel tree growth and batched rounds
Tree growth is sequential within a tree, but independent across trees given the local bootstrap and feature subsampling. We exploit this by training all trees level-wise: at a given depth, the server batches the active nodes across all trees into a single request, clients answer for the whole batch in one message, and the server updates all nodes independently. As a result, training $T$ trees to depth $M$ requires $O(M)$ synchronized rounds (rather than scaling with the total number of nodes). This batching is also compatible with client-side parallelism, since node computations are embarrassingly parallel across $\nu$.
4.3 FedForest Training Protocol
We summarize the complete server–client protocol for training a forest with $T$ trees of depth at most $M$ in Algorithm 1. Each node $\nu$ is identified by its path from the root and the tree index; each client can therefore compute the set of local samples reaching $\nu$ without sharing indices.
Algorithm 1 FedForest (Quantiles + Fisher for $H$)
1:Input: #trees $T$, depth $M$, subsample $r$, sketch size $B$.
@@ 4.3 FedForest Training Protocol
3:Init: Server selects client subsets $\mathcal{K}_{\tau}$ (size $\lceil rK\rceil$); Clients draw bootstrap samples.
4: Server initializes root nodes $\mathcal{A}=\{(\tau,\text{root})\}_{\tau=1}^{T}$.
5:for $\ell=0$ to M1M-1 do
@@
9:   Compute sketches $\mathbf{q}^{(j)}_{\nu,k}$ for continuous $X\in\mathcal{J}_{\nu}$.
10:   Compute local node summaries $\mathbf{s}_{\nu,k}$ (for $H$).
11:   Send sketches and summaries to Server.
12:  Server (Candidate Generation):
13:   For $X$: Aggregate sketches $\to$ approximate pooled CDF via Equation 4 $\to$ candidates $\mathcal{T}_{\nu}^{(X)}$.
@@
14:   For $H$: Compute site means from $\mathbf{s}_{\nu,k}$; sort site indices $\sigma$ by mean; generate splits $\mathcal{T}_{\nu}^{(H)}$.
15:   Broadcast candidates $\mathcal{T}_{\nu}=\mathcal{T}_{\nu}^{(X)}\cup\mathcal{T}_{\nu}^{(H)}$.
16:  Clients (Parallel): For all $\tau$ and $k\in\mathcal{K}_{\tau}$:
@@
17:   Compute left-child summaries $\mathbf{s}_{L,k}(j,t)$ for all candidates $(j,t)\in\mathcal{T}_{\nu}$.
18:   Send to Server.
19:  Server: Aggregates $\mathbf{s}_{L}=\sum_{k}\mathbf{s}_{L,k}$; evaluates impurity gain via Equation 5;
20:   Selects best split $(j^{*},t^{*})$ and updates $\mathcal{A}$.
21:end for

### v2

@@ 1 Introduction
Random Forests (RFs) (Breiman, 2001) are ensembles of CART decision trees (Breiman et al., 1984) that recursively partition the feature space into regions, referred to as nodes, by selecting splits that maximize an impurity reduction criterion (e.g., variance or Gini). Predictions are then produced by aggregating the outcomes associated with the terminal nodes, commonly called leaves. Despite the deep learning revolution, RFs remain among the state-of-the-art and most widely used predictors for tabular data (Fernández-Delgado et al., 2014; Grinsztajn et al., 2022; Kaggle, 2022; Uddin and Lu, 2024), while also offering practical advantages: they naturally handle both continuous and categorical features, require minimal hyperparameter tuning, and are substantially more computationally efficient than deep learning models (Grinsztajn et al., 2022). The strong empirical performance of RFs has been established in centralized settings, where data are assumed to be pooled together. However, this assumption is often unrealistic in many real-world applications, particularly in healthcare, where data are distributed across multiple institutions and cannot be centralized due to regulations, governance constraints, or competitive considerations (Rieke et al., 2020; Antunes et al., 2022; Nguyen et al., 2022; Xu et al., 2021). Federated learning (FL) addresses this challenge by enabling collaborative model training on decentralized data (Kairouz et al., 2021). In this work, we focus on the standard server–client architecture and on horizontal FL, where clients hold disjoint sets of records described by a common feature space.
@@
A recent survey by Wang and Gai (2024) identifies three key challenges in decentralized training of tree-based models: distributed computation of split statistics, limiting communication costs, and handling client heterogeneity.
@@
Although RFs enjoy a substantial theoretical foundation in the centralized setting (Biau and Scornet, 2016; Scornet and Hooker, 2026), analogous guarantees are currently lacking for their federated counterparts. To pave the way toward provably reliable federated forests, we study the fundamental question of whether a federated split-selection mechanism can recover, or closely approximate, the greedy split decisions made by centralized CART.
@@ 2.1 Centralized CART/RF Split Selection
The mixture-risk objective (1) is an ideal target. CART and Random Forests approximate it through greedy, node-wise optimization of an impurity criterion $I$ (e.g., variance for regression; Gini or entropy for classification) on a finite data sample. In the centralized setting, all data are pooled as $\mathcal{D}=\bigcup_{k=1}^{K}\mathcal{D}_{k}$. At a node $\nu\subseteq\mathcal{X}$ of a tree, let $\mathcal{S}_{\nu}\subseteq\mathcal{D}$ be the samples reaching $\nu$. CART selects a split threshold (j,t)(j,t)–that is, a split variable $j\in[d]$ and a cutpoint $t\in\mathbb{R}$–in two steps: (i) candidate generation, where for each feature $j$, the values $\{x_{i,j}\}_{i\in\mathcal{S}_{\nu}}$ are sorted, and candidates $t$ are defined as the midpoints between consecutive distinct values; and (ii) split evaluation, where each candidate $(j,t)$ is scored by its empirical impurity reduction
@@
with $\mathcal{S}_{L}=\{i\in\mathcal{S}_{\nu}:x_{i,j}\leq t\}$, $\mathcal{S}_{R}=\{i\in\mathcal{S}_{\nu}:x_{i,j}>t\}$, $n_{L}=|\mathcal{S}_{L}|$, $n_{R}=|\mathcal{S}_{R}|$, and $n_{\nu}=n_{L}+n_{R}$. $I(\mathcal{S}_{\nu})$ denotes the empirical impurity of the samples reaching node $\nu$; for instance, in a regression setting, it stands for the empirical variance of the outcomes of samples falling into $\mathcal{S}_{\nu}$. CART then selects the threshold $(j,t)$ that maximizes $\Delta I(j,t;\mathcal{S}_{\nu})$.
@@ 2.2 Heterogeneity Scenarios
Figure 1: Graphical model of heterogeneity regimes. Dashed arrows indicate possible dependencies between the client variable $H$, the covariates and/or outcome mechanism, depending on the regime.
General model. In full generality, the combined data admit the decomposition
@@
Assumption 2.2 (Covariate shift).
@@
Assumption 2.3 (Outcome shift).
$P(X\mid H)=P(X)$ while P(Y|X,H)P(Y|X{,H}) may depend on $H$.
@@ 3.1 Candidate Split Generation
Theorems 3.1 and 3.2 (proved in Appendix C.1) show that $B$ directly controls the rank resolution of the candidate grid: increasing $B$ makes the federated quantile thresholds closer to the centralized midpoint thresholds in terms of induced child assignments. For example, setting $B=100$ percentiles implies at most $1.5\%$ disagreement in child assignments while requiring only B1B-1 candidates per feature, independent of $n_{\nu}$. These results allow precise control of the trade-off between communication cost and candidate-threshold fidelity to the centralized approach.
Our quantile-based candidates offer several advantages to the fixed-width histograms proposed by Kalloori and Klingler (2022) for three reasons: (i) they adapt to local data density, ensuring critical split points are preserved as well as possible for a given budget $B$; (ii) they are more robust against outliers; and (iii) they ensure an efficient use of bins by enforcing a minimum sample fraction of $1/B$ within each bin (see experiment in Section F.3).
Top-$L$ feature shortlisting for homogeneous data. Under 2.1, local impurity gains concentrate around the pooled gain, so local rankings tend to agree with the global ranking. This enables a communication-efficient preselection step: each client $k$ computes for each feature $j$ its best local gain $M_{\nu,k}(j)=\max_{t}\ \Delta I(j,t;\mathcal{S}_{\nu,k})$ and sends only its top-$L$ features (with associated gains). The server merges these reports (e.g., taking the union of the local top-$L$ sets) to form a reduced feature set $\mathcal{L}_{\nu}$ on which it performs quantile sketching and split evaluation. A supporting guarantee for retaining the pooled-best feature under 2.1 is given in Appendix C.5.
@@ 3.2 Federated Impurity Reduction
Given the candidate set $\mathcal{T}_{\mathrm{fed}}$ obtained in Section 3.1, the remaining task is to evaluate the impurity reduction (2) for each candidate split $(j,t)$ without exchanging raw data. We first describe AvgImp, a natural evaluation rule for the homogeneous setting, and then introduce our main approach for the general heterogeneous setting.
Averaging local impurity gains under client homogeneity (AvgImp). Under 2.1, local split objectives are expected to align with the centralized CART objective. This motivates a simple federated evaluation rule based on averaging local impurity reductions. For each candidate split $(j,t)$, client $k$ computes its local gain $\Delta I(j,t;\mathcal{S}_{\nu,k})$, and the server forms the weighted average
The split is then selected by maximizing $\widehat{\Delta I}_{\mathrm{AvgImp}}$ over the $\mathcal{T}_{\mathrm{fed}}$. In practice, this can be combined with local Top-$L$ screening: each client reports only its best candidates according to local gains, and the server evaluates the averaged criterion on the resulting shortlist.
Theorem 3.3 (AvgImp approximation error on homogeneous data).
Under 2.1 and $\mathbb{E}[Y^{2}]<\infty$, we have:
While this approximation bound becomes uninformative in highly fragmented regimes where $K$ is comparable to $n$, it is meaningful in typical cross-silo federated learning scenarios, where $n\gg K$ (i.e., a small number of clients each with large datasets, such as hospitals with patient records).
However, the guarantee above relies on 2.1. In heterogeneous settings, we show that averaging local impurities does not generally recover the pooled impurity, as described in the following decomposition.
Theorem 3.4 (General impurity decomposition).
The impurity reduction for a given candidate decomposes as
$\Delta I(j,t;\mathcal{S}_{\nu})=\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}\Delta I(j,t;\mathcal{S}_{\nu,k})+\Delta\mathcal{E}(j,t;\nu),$ (5)
where $\Delta\mathcal{E}(j,t;\nu)$ is a split-dependent heterogeneity correction, given in Section C.3. It corresponds to the change in the Jensen-gap term $\mathcal{E}(\nu)$ caused by splitting the within-node mixture of client distributions. For standard strictly concave impurity functions (e.g., Gini index, entropy, and variance), we have $\mathcal{E}(\nu)\geq 0$, with $\mathcal{E}(\nu)=0$ if and only if
that is, the within-node empirical outcome means, or class probabilities, coincide across all clients.
Thus, when clients hold heterogeneous distributions, split decisions based on AvgImp will deviate from those produced by a centralized CART procedure. This discrepancy motivates an alternative evaluation rule that remains exact even under heterogeneous data.
Exact impurity evaluation via additive summaries. Our main evaluation rule relies on the novel observation that standard CART impurities admit additive sufficient statistics, as shown in Theorem 3.5. These statistics can be aggregated across clients to recover the centralized impurity reduction exactly for any candidate split.
Theorem 3.5 (Exact centralized evaluation via additive summaries).
For standard CART criteria, there exists a deterministic map $\Psi$ and per-client summaries $\mathbf{s}_{\nu,k}$ such that the pooled impurity reduction can be written as
@@
with aggregated statistics $\mathbf{s}_{m}=\sum_{k=1}^{K}\mathbf{s}_{m,k}$ for $m\in\{\nu,L,R\}$.
In practice, for each candidate $(j,t)$, client $k$ sends only its left-child summary $\mathbf{s}_{L,k}$. The server then aggregates these summaries into $\mathbf{s}_{L}$, recovers $\mathbf{s}_{R}=\mathbf{s}_{\nu}-\mathbf{s}_{L}$, and evaluates (6). This procedure exactly replicates centralized split evaluation, irrespective of covariate or outcome shift.
Examples. For $C$-class classification, $\mathbf{s}_{\nu}=(n_{\nu,1},\dots,n_{\nu,C})$ and $\Psi_{\mathrm{Gini}}(\mathbf{s}_{\nu})=1-\sum_{c=1}^{C}(n_{\nu,c}/n_{\nu})^{2},$ so each client sends $C$ scalars per candidate. For regression, $\mathbf{s}_{\nu}=(n_{\nu},S_{\nu},Q_{\nu})$, with $S_{\nu}=\sum y_{i}$ and $Q_{\nu}=\sum y_{i}^{2}$, and $\Psi_{\mathrm{Var}}(\mathbf{s}_{\nu})=Q_{\nu}/n_{\nu}-(S_{\nu}/n_{\nu})^{2},$ so each client sends three scalars per candidate.
In summary, AvgImp provides a natural evaluation rule under homogeneous client distributions, but its guarantees do not extend to violations of 2.1. The additive-summary rule in Theorem 3.5 is therefore our primary evaluation rule when exact centralized split selection is required.
Outcome shift: splitting on $H$ with no extra communication. When 2.3 holds (which can be tested in a federated manner, see Section F.4), the target rule is $f(x,h)=\mathbb{E}[Y\mid X=x,H=h]$, and optimal trees may require splits on $H$. Let $\mathcal{K}_{\nu}=\{k:\ n_{\nu,k}>0\}$ be the set of active clients at node $\nu$. Once the server has collected the per-client node summaries $\{\mathbf{s}_{\nu,k}\}_{k\in\mathcal{K}_{\nu}}$—already needed to evaluate feature-based splits—it can also evaluate splits on $H$ at no additional communication cost, i.e., without requesting any extra messages from clients. Indeed, any partition $\mathcal{K}_{\nu}=\mathcal{K}_{L}\cup\mathcal{K}_{R}$ induces $\mathbf{s}_{\nu,L}=\sum_{k\in\mathcal{K}_{L}}\mathbf{s}_{\nu,k}$ and $\mathbf{s}_{\nu,R}=\sum_{k\in\mathcal{K}_{R}}\mathbf{s}_{\nu,k},$ and the corresponding gain follows directly from (6) using the same map $\Psi$.
@@ 4 Federated Forests
We now show how to lift the federated CART primitives from Section 3—candidate generation and summary-based split evaluation—into a scalable Random Forest training protocol, which we call FedForest. The protocol, described in Algorithm 1, rests on three key components.
Client-stratified bootstrap. For each tree, client $k$ draws (with replacement) a bootstrap sample of size $n_{k}$ from its local data $\mathcal{D}_{k}$, and the tree is trained on the virtual union of these resamples across clients (Alg. 1, Line 4). This preserves the pooled mixture weights $\rho_{k}=n_{k}/n$ in every tree. In contrast, a naive pooled bootstrap can under-represent—or entirely omit—small clients.
Feature subsampling. As in standard Random Forests, we introduce tree-level randomness via feature subsampling to promote diversity across trees. For each active node, the server samples a subset of features $\mathcal{J}_{\nu}$ to evaluate (Alg. 1, Line 6). In the federated setting, this also reduces communication, as fewer feature–threshold pairs need to be evaluated and transmitted by each client. It can further be combined with Top-$L$ shortlisting to achieve additional savings.
Level-wise batched training. Tree growth is sequential within each tree but independent across trees. We leverage this by training all trees level-wise (Alg. 1, Lines 5–18): at each depth, the server batches active nodes across all trees into a single request. Clients answer for the whole batch in parallel (Alg. 1, Lines 7–11 and 15–17), and the server updates all nodes independently. As a result, training $T$ trees of depth $M$ requires only $O(M)$ synchronized rounds, rather than scaling with the total number of nodes.
Algorithm 1 FedForest
1:Input: #trees $T$, depth $M$, client subsampling rate $r$, sketch size $B$.
@@
3:Init: Server initializes root nodes $\mathcal{A}=\{(\tau,\text{root})\}_{\tau=1}^{T}$.
4: For each $\tau$, server selects client subsets $\mathcal{K}_{\tau}$ (size $\lceil rK\rceil$); Clients draw bootstrap samples.
5:for depth level $\ell=0$ to M1M-1 do
@@
9:   Compute sketches $\mathbf{q}^{(j)}_{\nu,k}$ for continuous $X\in\mathcal{J}_{\nu}$ and local node summaries $\mathbf{s}_{\nu,k}$ (for $H$).
10:   Send sketches and summaries to Server.
11:  Server (Candidate Generation):
12:   For $X$: Aggregate sketches $\to$ approximate global CDF via Eq. 4 $\to$ candidates $\mathcal{T}_{\nu}^{(X)}$.
@@
13:   For $H$: Compute site means from $\mathbf{s}_{\nu,k}$; sort site indices by mean; generate splits $\mathcal{T}_{\nu}^{(H)}$.
14:   Broadcast candidates $\mathcal{T}_{\nu}=\mathcal{T}_{\nu}^{(X)}\cup\mathcal{T}_{\nu}^{(H)}$.
15:  Clients (Parallel): For all $\tau$ and $k\in\mathcal{K}_{\tau}$:
@@
16:   Compute left-child summaries $\mathbf{s}_{L,k}(j,t)$ for all candidates $(j,t)\in\mathcal{T}_{\nu}$ and send to Server.
17:  Server: Aggregates $\mathbf{s}_{L}=\sum_{k}\mathbf{s}_{L,k}$; evaluates impurity gain via Equation 6;
18:   Selects best split $(j^{*},t^{*})$ and updates $\mathcal{A}$.